# Graph quandles: Generalized Cayley graphs of racks and right quasigroups

### v1

@@ 1. Introduction
Consider the class of all deterministic, source-complete labeled digraphs. The subclass whose elements are codeterministic and target-complete (resp. codeterministic, target-complete) is precisely the class $\mathcal{Q}$ of labeled Cayley digraphs of right quasigroups (resp. right-cancellative magmas, right-divisible magmas). (See Theorem 7.12.)
Moreover, there exist graph theoretic-conditions on $\mathcal{Q}$ that restrict it to the subclass of labeled Cayley digraphs of racks, and similarly for quandles, involutory racks, and kei. (See Theorem 7.14.)
@@ 1.1.3. Characterization of labeled Cayley digraphs
[…]
@@
Caucal answered Problem 1.8 for left quasigroups, quasigroups [caucal], semigroups, and various classes of monoids, including groups [caucal2]. [^2] Since the right-multiplication maps $R_{v}$ of left quasigroups are not necessarily permutations, the answers to Problem 1.8 for left and right quasigroups are distinct. Chishwashwa et al. addressed a similar question for vertex-labeled Cayley digraphs of unital, fixed point-free right quasigroups [groupoid2]. In this paper, we answer Problem 1.8 for various classes of racks and right quasigroups; see Theorems 7.12 and 7.14.
@@ 1.3. Notation
Acknowledgments
I thank Sam Raskin for suggesting Proposition 2.11 while advising a previous paper. I thank Casey Tompkins for introducing me to algebraic graph theory, and I thank Jose Ceniceros and Peyton Wood for introducing me to quandle theory.
@@ 2.1. Magmas
A magma or groupoid is a pair $(V,R)$, denoted by $V_{R}$, where $V$ is a set and $R$ is a mapping from $V$ to the set of functions from $V$ to $V$ [^3] Although the notation $V_{R}$ is nonstandard, we use it because Bardakov formulated his open problems with it [bardakov]. For all $v\in V$, we call the map $R_{v}:=R(v)$ a right-multiplication map or right-translation map.
@@
The addition maps $R_{y}(x):=x+y$ define a right quasigroup structure on the nonzero rational numbers $\mathbb{Q}\setminus\{0\}$ that is not a group structure. With respect to these maps, the positive rational numbers $\mathbb{Q}^{+}$ are a submagma of $\mathbb{Q}\setminus\{0\}$ but not a right quasigroup.
@@ 3.2. Marked graphs
If $n=2$, it is easy to see that of the eight possible markings $R:V\to S_{n}$, exactly two are rack structures. Indeed, if $\ell_{1},\ell_{2}\in V$ are the two leaves and $v\in V$ is the central vertex, then Equation (2.1) forces $R_{\ell_{1}}=R_{\ell_{2}}$ and $R_{v}=\operatorname{id}_{V}$. For an example of a marking of $K_{1,2}$ that does not realize a rack, see Example 4.1.
@@ 3.4. Schreier graphs
Given a right quasigroup $V_{R}$ and a subset $S\subseteq V$, let $G:=\operatorname{RMlt}V_{R}$ and ${T:=R(S)}$. Then
@@ 5. From marked graphs to racks
$\mu_{\mathrm{rack}}(P_{n})=2^{n-1},\quad\quad\mu_{\mathrm{qnd}}(P_{n})=2^{n-2}.$
@@
Let $\ell_{1},\ell_{2}\in V(P_{n})$ be the leaves. Then the nonidentity element of $\operatorname{Aut}P_{n}\cong\mathbb{Z}/2\mathbb{Z}$ swaps $\ell_{1}$ and $\ell_{2}$ and fixes all other vertices. It is easy to verify that a map $R:V(P_{n})\to\operatorname{Conj}(\operatorname{Aut}P_{n})$ is a magma homomorphism if and only if $R_{\ell_{1}}=R_{\ell_{2}}$. In this case, $R$ is a quandle structure if and only if $R_{\ell_{1}}=\operatorname{id}_{V(P_{n})}$. Hence, the claim follows from Theorem 5.1. ∎
@@
When $n=3$, five markings of the cycle graph $C_{3}$ realize quandles. In particular, let $V_{R}$ be the quandle from Example 4.3. The full Cayley graph $\Gamma(V_{R})\cong C_{3}$ depicted in Figure 3, marked by the quandle structure $R:V\to S_{3}=\operatorname{Aut}C_{3}$, realizes $V_{R}$. In the following section, we generalize this by showing that all racks are realized by their full Cayley (di)graphs.
@@ 7.1.2. Labeled digraphs
Let $\mathcal{D}$ denote the class of deterministic, source-complete labeled digraphs. Let $\mathcal{Q}$ denote the subclass of $\mathcal{D}$ whose elements are also codeterministic and target-complete.
@@ 7.3. First results
The following answers Problem 1.8 for right-cancellative magmas, right-divisible magmas, and right quasigroups.
@@ 7.4. Specialization to racks
  • We say that $\Gamma$ satisfies the rack condition if for all vertices $v\in V$, if $(v,\ell_{1},w_{1}),(v,\ell_{2},w_{2})\in E$ are edges of $\Gamma$, then
@@
  • We say that $\Gamma$ satisfies the label-idempotence condition if $R_{\ell}(\ell)=\ell$ for all $\ell\in L$; that is, each vertex $\ell$ contained in the labeling set $L$ has a loop labeled by $\ell$.
  • We say that $\Gamma$ satisfies the label-involution condition if $R^{2}_{\ell}(v)=v$ for all $v\in V$ and $\ell\in L$; that is, the edge set $E$ can be partitioned into loops and cycles of length 2 having the form
@@
[…]
    Figure 7. The label-involution condition from Definition 7.13 states that subgraphs of the forms on the left and the right partition the edge set.
@@ 7.4.1. Result
    $\Gamma$ is the labeled Cayley digraph of a rack if and only if $\Gamma$ lies in $\mathcal{Q}$ and satisfies the rack condition.
@@
    $\Gamma$ is the labeled Cayley digraph of a quandle if and only if $\Gamma$ lies in $\mathcal{Q}$ and satisfies the rack and label-idempotence conditions.
@@
    $\Gamma$ is the labeled Cayley digraph of an involutory right quasigroup if and only if $\Gamma$ lies in $\mathcal{Q}$ and satisfies the label-involution condition.
@@
(1): First, suppose that $V^{\Gamma}_{R}$ is a rack; we show that $\Gamma$ satisfies the rack condition. For all vertices $v\in V$ and edges ${(v,\ell_{i},w_{i})\in E}$, we have $w_{i}=R_{\ell_{i}}(v)$. It follows from Equation (2.1) that
@@
Since $v\in V$ and ${(v,\ell_{i},w_{i})\in E}$ were arbitrary, $\Gamma$ satisfies the rack condition.
Conversely, suppose that $\Gamma$ satisfies the rack condition; we show that $V^{\Gamma}_{R}$ is a rack. We have to verify that
@@
for all $\ell_{1},\ell_{2},v\in V$. If $\ell_{1}\notin L$ or $\ell_{2}\notin L$, we are done. Otherwise, let $w_{i}:=R_{\ell_{i}}(v)$ for $i=1,2$, so $(v,\ell_{i},w_{i})\in E$. Since $\Gamma$ satisfies the rack property,
$R_{\ell_{1}}R_{\ell_{2}}(v)=R_{\ell_{1}}(w_{2})=R_{R_{\ell_{1}}(\ell_{2})}(w_{1})=R_{R_{\ell_{1}}(\ell_{2})}R_{\ell_{1}}(v),$
as desired.
@@ 8. Open questions
Define and characterize Cayley (di)graphs of various classes of racks equipped with additional structure (e.g., generalized Legendrian racks [ta], multi-virtual quandles [virtual], symmetric racks [ta2]).

### v3

@@ 1. Introduction
Let $\mathcal{Q}$ be the class of all labeled digraphs that are deterministic, source-complete, codeterministic, and target-complete. Then $\mathcal{Q}$ is precisely the class of labeled Cayley digraphs of right quasigroups. (See Theorem 7.12.)
Moreover, there exist graph theoretic-conditions on $\mathcal{Q}$ that restrict it to the subclass of labeled Cayley digraphs of racks, and similarly for quandles, involutory racks, and kei. (See Theorem 7.14 and Corollary 7.15.)
@@ 1.1.3. Characterization of labeled Cayley digraphs
Although this paper is the first to study labeled Cayley (di)graphs of right quasigroups in general, various authors have studied Cayley graphs of various classes of right quasigroups appearing in combinatorics, algebraic topology, and knot theory. For example, full Cayley digraphs of unital, fixed point-free right quasigroups have various applications in network theory [groupoid1], and full Cayley graphs of right quasigroups have applications in categorical covering theory [categories].
@@
Caucal answered Problem 1.8 for left quasigroups, quasigroups [caucal], semigroups, and various classes of monoids, including groups [caucal2]. [^2] Since the right-multiplication maps $R_{v}$ of left quasigroups are not necessarily permutations, the answers to Problem 1.8 for left and right quasigroups are distinct. Chishwashwa et al. addressed a similar question for vertex-labeled Cayley digraphs of unital, fixed point-free right quasigroups [groupoid1]. In this paper, we answer Problem 1.8 for various classes of racks and right quasigroups; see Theorems 7.12 and 7.14 and Corollary 7.15.
@@ 2.1. Magmas
A magma or groupoid is a pair $(V,R)$, denoted by $V_{R}$, where $V$ is a set and $R$ is a mapping from $V$ to the set of functions from $V$ to VV. [^3] Although the notation $V_{R}$ is nonstandard, we use it because Bardakov formulated his open problems with it [bardakov]. For all $v\in V$, we call the map $R_{v}:=R(v)$ a right-multiplication map or right-translation map.
@@
If we take $G:=\mathbb{Q}$ to be the rational numbers in Example 2.2, then the positive rational numbers $\mathbb{Q}^{+}$ are a submagma of $\mathbb{Q}$ but not a right quasigroup.
@@ 3.2. Marked graphs
If $n=2$, it is easy to see that of the eight possible markings $R:V\to S_{n}$, exactly four are rack structures. Indeed, if $\ell_{1},\ell_{2}\in V$ are the two leaves and $v\in V$ is the central vertex, then Equation (2.1) forces $R_{\ell_{1}}=R_{\ell_{2}}$. For an example of a marking of $K_{1,2}$ that does not realize a rack, see Example 4.1.
@@ 3.4. Schreier graphs
Given a right quasigroup $V_{R}$ and a subset $S\subseteq V$, let $G:=\operatorname{RMlt}V_{R}$ and T:=R(S){T:=R(S)}. Then
@@ 5. From marked graphs to racks
$\mu_{\mathrm{rack}}(P_{n})=\begin{cases}2^{k}&\text{if }n=2k,\\ 2^{k+1}&\text{if }n=2k+1,\end{cases}\qquad\mu_{\mathrm{qnd}}(P_{n})=\begin{cases}1&\text{if }n=2k,\\ 2&\text{if }n=2k+1.\end{cases}$
@@
Let $V(P_{n})=[n]$ be the vertices of $P_{n}$ in order, that is, the edges of $P_{n}$ are $(i,i+1)$ for all $1\leq i\leq n-1$. Then the nonidentity element $\sigma$ of $\operatorname{Aut}P_{n}\cong\mathbb{Z}/2\mathbb{Z}$ swaps $i$ and n+1in+1-i for all $1\leq i\leq\lfloor n/2\rfloor$. Therefore, a function $R:V(P_{n})\to\operatorname{Conj}(\operatorname{Aut}P_{n})$ is a magma homomorphism if and only if $R_{i}=R_{n+1-i}$ for all $1\leq i\leq n-1$. Since $\sigma$ has no fixed points if $n=2k$ and a unique fixed point when $n=2k+1$, the claim follows from Theorem 5.1. ∎
@@
When $n=3$, five markings of the cycle graph $C_{3}$ realize quandles. In particular, let $V_{R}$ be the quandle from Example 4.2. The full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)\cong C_{3}$ depicted in Figure 2, marked by the quandle structure $R:V\to S_{3}=\operatorname{Aut}C_{3}$, realizes $V_{R}$. In the following section, we generalize this by showing that all racks are realized by their full Cayley (di)graphs.
@@ 7.1.2. Labeled digraphs
Let $\mathcal{D}$ denote the class of deterministic, source-complete labeled digraphs.
@@ 7.3. First results
Henceforth, let $\mathcal{Q}$ denote the subclass of $\mathcal{D}$ whose elements are also codeterministic and target-complete. The following answers Problem 1.8 for right-cancellative magmas, right-divisible magmas, and right quasigroups.
@@ 7.4. Specialization to racks
  • We say that $\Gamma$ satisfies the first rack condition if for all $v\in V$, if $(v,\ell_{1},w_{1}),(v,\ell_{2},w_{2})\in E$,
@@
  • We say that $\Gamma$ satisfies the second rack condition if for all edges $(v,\ell,w)$ and non-label vertices $x\in V\setminus L$ such that $R_{\ell}(x)\in L$ is a label, then $w$ has a loop labeled by $R_{\ell}(x)$; i.e.,
    $(w,R_{\ell}(x),w)\in E.$
    See Figure 7 for a visualization.
    $R_{\ell}(x)\ellvvwwww$
    Figure 7. The second rack condition from Definition 7.13 states that for all subgraphs of the form on the left and non-label vertices $x\in V\setminus L$, if $R_{\ell}(x)\in L$ is a label, then there exists a subgraph of the form on the right.
  • We say that $\Gamma$ is label-idempotent if $R_{\ell}(\ell)=\ell$ for all $\ell\in L$; that is, each vertex $\ell$ contained in the labeling set $L$ has a loop $(\ell,\ell,\ell)\in E$ labeled by $\ell$.
  • We say that $\Gamma$ satisfies the label-involutory if $R^{2}_{\ell}(v)=v$ for all $v\in V$ and $\ell\in L$; that is, the edge set $E$ can be partitioned into loops and cycles of length 2 having the form
@@
    $\ell\ell\ellvvvvww$
    Figure 8. The label-involutory condition from Definition 7.13 states that subgraphs of the forms on the left and the right partition the edge set.
@@ 7.4.1. Result
    $\Gamma$ is the labeled Cayley digraph of a rack if and only if $\Gamma$ lies in $\mathcal{Q}$ and satisfies the two rack conditions.
@@
    $\Gamma$ is the labeled Cayley digraph of a quandle if and only if $\Gamma$ lies in $\mathcal{Q}$, satisfies the two rack conditions, and is label-idempotent.
@@
    $\Gamma$ is the labeled Cayley digraph of an involutory right quasigroup if and only if $\Gamma$ is a label-involutory element of $\mathcal{Q}$.
@@
(1): First, suppose that $V^{\Gamma}_{R}$ is a rack; we show that $\Gamma$ satisfies the rack conditions. For all vertices $v\in V$ and edges $(v,\ell_{i},w_{i})\in E$}, we have $w_{i}=R_{\ell_{i}}(v)$. It follows from Equation (2.1) that
@@
Hence, $\Gamma$ satisfies the first rack condition. Next, let $(v,\ell,w)\in E$, and let $x\in V\setminus L$ satisfy $R_{\ell}(x)\in L$. Since $R_{x}=\operatorname{id}_{V}$ and $R_{\ell}(v)=w$, Equation (2.1) yields
$R_{R_{\ell}(x)}(w)=R_{R_{\ell}(x)}R_{\ell}(v)=R_{\ell}R_{x}(v)=R_{\ell}(v)=w.$
Since $\Gamma=\Gamma_{\mathrm{lab}}(V^{\Gamma}_{R},L)$, it follows that $(w,R_{\ell}(x),w)\in E$, so $\Gamma$ satisfies the second rack condition.
Conversely, suppose that $\Gamma\in\mathcal{Q}$ satisfies the two rack conditions; we show that $V^{\Gamma}_{R}$ is a rack. We have to verify that
@@
for all $\ell_{1},\ell_{2},v\in V$. If $\ell_{1}\notin L$, we are done. Next, suppose that $\ell_{1}\in L$ and $\ell_{2}\notin L$. If $R_{\ell_{1}}(\ell_{2})\notin L$, we are done. Otherwise, applying the second rack condition to the edge $(v,\ell_{1},R_{\ell_{1}}(v))$ yields the desired equality. Finally, if $\ell_{1},\ell_{2}\in L$, then $(v,\ell_{i},w_{i})\in E$ with $w_{i}:=R_{\ell_{i}}(v)$. Since $\Gamma$ satisfies the first rack property,
$R_{\ell_{1}}R_{\ell_{2}}(v)=R_{\ell_{1}}(w_{2})=R_{R_{\ell_{1}}(\ell_{2})}(w_{1})=R_{R_{\ell_{1}}(\ell_{2})}R_{\ell_{1}}(v).$
@@
Corollary 7.15.
Let $\mathcal{I}$ be the subclass of $\mathcal{Q}$ whose elements are label-involutory and satisfy the two rack conditions. Then $\mathcal{I}$ (resp. the subclass of $\mathcal{I}$ whose elements are label-idempotent) is precisely the class of labeled Cayley digraphs of involutory racks (resp. kei). ∎
@@ 8. Open questions
Define and characterize Cayley (di)graphs of classes of racks equipped with extra structure (e.g., generalized Legendrian racks [ta], multi-virtual quandles [virtual], symmetric racks [ta2]).